OpenStudy (anonymous):

If kelly received 1/3 more votes than mike in a student election, which of the following could have been the total number of votes cast for the two candidates? a)12 b)13 c)14 d)15 e)16 thanks!

OpenStudy (anonymous):

15 votes

OpenStudy (anonymous):

Kelly = 9, Mike = 6 , Total = 15

OpenStudy (anonymous):

wrong. the answer is c.

OpenStudy (anonymous):

plz explain @sania22

OpenStudy (anonymous):

i figured it out. if kelly=1/3M+M=4/3M. the total is M+(4/3M)= 7/3M. T=7/3M. then M has to be a whole number, so M=3/7T and has to be a multiple of 7, the only multiple would be c)14.

OpenStudy (anonymous):

ya u r right